# Phase space analysis of CCDM cosmologies

### v1

@@ III.5 Model: $\Gamma=\Gamma_{0}H^{-2}$
In this series the last model is $\Gamma=\Gamma_{0}H^{-2}$ where $\Gamma_{0}$ is a constant with dimension equal to the dimension of the cube of the Hubble rate.
III.5.1 One fluid system
In the single fluid system, one dimensional autonomous system becomes,
$z^{\prime}=\frac{3}{2}z\left(1-z-\frac{\mu z^{3}}{3(1-z)^{2}}\right),$ (39)
which has the singularity at $z=1$ and $\mu=\frac{\Gamma_{0}}{H_{0}^{3}}$ is a parameter which always takes positive value. Removing this singularity by multiplying the positive factor $(1-z)^{2}$, the above system reduced to its regularized form:
$z^{\prime}=\frac{3}{2}z\left((1-z)^{3}-\frac{\mu z^{3}}{3}\right).$ (40)
Since, the decelerating parameter $q=\frac{1}{2}\left[1-\frac{\mu z^{3}}{(1-z)^{3}}\right]$ at present time (i.e. $z=1/2$) is $q=\frac{1-\mu}{2}$, hence, for the present time to have an accelerating phase of the universe, we should have $\mu>1$. This system has two critical points, namely $E^{\prime}_{0}$ and $E^{\prime}_{1}$ which are shown in Table 5. In $E^{\prime}_{0}$ and $E^{\prime}_{1}$, the first component corresponds to DM density parameter ($\Omega_{\rm dm}$), which always equals to $1$ and the second component stands for the $z$ variable. Both the critical points always represent matter dominated phase, but $E^{\prime}_{0}$ is an unstable point showing deceleration and $E^{\prime}_{1}$ gives acceleration with stable behavior. The stability behavior of the critical points are represented in the upper left plot of Fig. 9 by the direction of the vector field. The cosmological features are identical with the one-fluid system of the Models III.1 and III.4.
III.5.2 Two fluid system
@@
where $\mu=\frac{\Gamma_{0}}{H_{0}^{3}}$ is a dimensionless parameter adopting positive values. But as before, there is a singularity in the system at $z=1$ and we shall take aside this issue by regularization procedure using the positive factor $(1-z)^{3}$ and this leads to the following system
$x^{\prime}=$ $x(1-x)\left(3w(1-z)^{3}+\mu z^{3}\right),$ (42a)
$z^{\prime}=$ $\frac{3}{2}z(1-z)\left[\left(1+w(1-x)\right)(1-z)^{3}-\frac{\mu xz^{3}}{3}\right].$ (42b)
The Friedmann equation (11) and the dimensionless variables $x,~{}z$ defined in (24) gives our physical region which is again a unit square and we have named it as $R_{5}=[0,1]^{2}$. Here one can easily conclude that $R_{5}$ is a positively invariant set under the dynamical system (42) because boundaries of $R_{5}$ are invariant manifolds of (42). Here, $q=\frac{1}{2}\left[1+3w(1-x)-\frac{\mu xz^{3}}{(1-z)^{3}}\right]$. Therefore, at present time (i.e. $z=1/2$) for the accelerating expansion of the universe, the physical trajectories must follow the condition $(1-\mu)x+(1+3w)(1-x)<0$. The cosmological features that are obtained from the above system are briefly given in the Table 5. Now, we split the equation of state parameter in various parts and analyse the system (42) in each part of $w$, which have been described below:
One fluid system
Critical point Existence Stability $\Omega_{\rm dm}$ $q$ Acceleration
$E^{\prime}_{0}(1,0)$ Always Always Unstable 1 $\frac{1}{2}$ No
$E^{\prime}_{1}\left(1,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$ $\mu>0$ Always Stable 1 $-1$ Yes
Two fluid system
Critical point Existence Eigenvalue Stability $\Omega_{f}$ $\Omega_{\rm dm}$ $q$ Acceleration
$E_{0}(0,0)$ Always $\left(3w,\frac{3}{2}(1+w)\right)$ Stable if $w<-1$; 1 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Saddle if $-1<w<0$;
@@
$E_{2}(0,1)$ Always $(\mu,0)$ Non-hyperbolic Saddle if $w>-1$; 1 0 Undefined Undetermined
Unstable if $w<-1$
$E_{3}(1,1)$ Always $\left(-\mu,\frac{\mu}{2}\right)$ Always Saddle 0 1 $\infty$- ∞ Yes
$E_{4}\left(1,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$ $\mu>0$ $\left(-\frac{\mu(1+w)}{\left(1+\sqrt[3]{\frac{\mu}{3}}\right)^{3}},-\frac{3\mu}{2\left(1+\sqrt[3]{\frac{\mu}{3}}\right)^{3}}\right)$ Stable if $w>-1$ 0 1 $-1$ Yes
Saddle if $w<-1$
$E_{5}\left(0,z_{c}\right)$ $w=-1$ $\left(\mu z_{c}^{3}-3\left(1-z_{c}\right)^{3},0\right)$ Stable if $z_{c}<\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$ and $1$ 00 $-1$ Yes
@@
$E_{6}\left(x_{c},\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$ $w=-1$ $\left(0,-\frac{3\mu x_{c}}{2\left(1+\sqrt[3]{\frac{\mu}{3}}\right)^{3}}\right)$ Stable $1-x_{c}$ $x_{c}$ $-1$ Yes
@@
Table 5: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (40) and (42) with the matter creation rate $\Gamma=\Gamma_{0}H^{-2}$.
@@
Figure 9: Description of the phase space structure of the one-fluid and two-fluid system for the matter creation rate $\Gamma=\Gamma_{0}H^{-2}$. Upper Left Plot: This is the vector plot of the one dimensional system (40) considering $\mu=3$. It is also noted that for other positive values of $\mu$, we can find similar type of plot. Upper Right Plot: The phase plot of the system (42) when we have assumed $w=0.5$ and $\mu=2$. For other values of $w>0$ and $\mu>0$, we can also obtain similar phase space structure. Lower Left Plot: The phase space of the system (42) when the equation of state $w$ takes the value 00. Here we use $\mu=2$ but any positive value of $\mu$ gives same type of phase portrait. Lower Right Plot: The phase plot of the system (42) considering $w=-0.3$ and $\mu=3$. Also, we can get similar type of graphics for any positive value of $\mu$ and negative value of $w$ in the interval $(-1,0)$. Here the green region corresponds to the decelerating phase ($q>0$), the cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase (i.e. $q<-1$). The black dotted curve separating the cyan and yellow regions corresponds to $q=-1$.
Figure 10: This figure depicts the phase space structure of the two-fluid system for the matter creation rate $\Gamma=\Gamma_{0}H^{-2}$ but here we consider the case when $w\leq-1$. Left Plot: The phase space of the system (42) when the equation of state $w$ takes the value $-1$. Here we use $\mu=2$ but any positive value of $\mu$ gives similar type of phase portrait. Right Plot: The phase plot of the system (42) when we assume $w=-2$ and $\mu=2$. For other values of $w<-1$ and $\mu>0$, we can also obtain similar graphics. Here the green region corresponds to the decelerating phase ($q>0$), the cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase (i.e. $q<-1$). The black dotted curve separating the cyan and yellow regions (in the left graph this curve is not clear because of the critical line) corresponds to $q=-1$.
  1. 1.
@@
    When $w>0$: For this part, the dynamical system (42) allows five critical points $E_{0},~{}E_{1},~{}E_{2},~{}E_{3}$ and $E_{4}$. Here, the accelerating phase includes the points $E_{3},~{}E_{4}$ and the decelerating phase contains the critical points $E_{0},~{}E_{1}$. Again, $E_{1},~{}E_{3},~{}E_{4}$ correspond to matter dominated solutions, while $E_{0},~{}E_{2}$ represent completely second fluid dominated points. One can easily conclude about stability from the direction of the vector field on the boundary of the phase space and signs of the eigenvalues. Therefore, $E_{0}$ is unstable, $E_{4}$ is the only globally stable point and $E_{1},~{}E_{2},~{}E_{3}$ are saddle by nature. Note that, $E_{4}$ is the DM dominated late time accelerating globally stable point. The upper right plot of Fig. 9 gives clear picture of the qualitative behavior of critical points. The inherent features of this plot are similar to the case $w>0$ of the Models III.1 and III.4.
  2. 2.
@@
    When $w=0$: When the equation of state $w$ of the second fluid vanishes, we attain three isolated critical points $E_{2}(0,1),~{}E_{3}(1,1)$ and $E_{4}\left(1,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$ along with one critical lines, namely, $E_{7}(x_{c},0)$. Now $x$ is always increasing and, $z$ is increasing and decreasing depending on $z$ variable chosen from $\left(0,\frac{1}{1+\sqrt[3]{\frac{\mu x}{3}}}\right)$ and $\left(\frac{1}{1+\sqrt[3]{\frac{\mu x}{3}}},1\right)$ respectively. Therefore $E_{2},~{}E_{3}$ are saddle, $E_{7}$ is unstable and $E_{4}$ is a globally stable critical point, which are obtained in the lower left plot of Fig. 9. Once again $E_{4}$ corresponds to DM dominated late time accelerated globally stable point. In this case, the cosmological characteristics are completely identical with the case $w=0$ of the Models III.1 and III.4.
  3. 3.
@@
    When $-\frac{1}{3}\leq w<0$: Similar to the case where $w>0$, in this part of $w$, we acquire five isolated critical points which are depicted in the Table 5. From this table and the flow on the unit square, we can say that $E_{1}$ is unstable i.e. the DM dominated point $E_{1}$ leaves its decelerating phase, $E_{0},~{}E_{2},~{}E_{3}$ are saddle by behavior and once again, $E_{4}$ is the DM dominated accelerating globally stable point. Therefore, the universe ends with DM dominated accelerating phase. The orbits in the phase space are highlighted in the lower right plot of Fig. 9, providing the results which replicate the case $-1/3<w<0$ for the Models III.1 and III.4.
  4. 4.
@@
    When $-1<w<-\frac{1}{3}$: This part of $w$ is analogous to the above case where $w$ lies in the interval $\left[-\frac{1}{3},0\right)$, except decelerating parameter at the critical point $E_{0}$ which belongs to the accelerating region in the phase space. Therefore, the phase plot is similar to the lower right plot of Fig. 9.
  5. 5.
@@
    When $w=-1$: If the equation of state of non-matter creating fluid $w$ coincides with cosmological constant, we can obtain two isolated critical points $E_{1},~{}E_{3}$ which are mentioned in Table 5 and two critical lines, namely, $E_{5}(0,z_{c}),~{}E_{6}\left(x_{c},\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$. It is clear that $x^{\prime}$ is positive and negative respectively if we choose $z$ from the interval $\left(\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}},1\right]$ and the interval [$0,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$. Again $z$ is increasing and decreasing below and above the line $z=\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$ respectively. In the phase space below the line $z=\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$, the equation of the separatrix is $z=\sqrt[3]{\frac{3}{\mu}}(1-z)\sqrt{1-x}$ which connects the critical points $E_{1}$ and $\left(0,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$. As a result, $E_{1}$ is unstable, $E_{3}$ is saddle, $E_{6}$ is stable and, $E_{5}$ is stable when $z_{c}<\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$ and unstable when $z_{c}>\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$. The left plot of Fig. 10 depicts qualitative features of the critical points. Here, $E_{5}$, lying in the accelerating phase, corresponds to completely DE domination and $E_{6}$, belonging also to the accelerating phase, represents the points where DM, DE coexist except the end points. Clearly, $E_{6}$ can solve cosmic coincidence problem. The key findings of this case are equivalent to the Models III.1 and III.4 with $w=-1$.
  6. 6.
@@
    When $w<-1$: Here, one can get two unstable critical points $\left(E_{1},E_{2}\right)$, two saddle type critical points $\left(E_{3},E_{4}\right)$ and $E_{0}$ which is the only globally stable critical point. From Table 5, it is clear that, $E_{1}$ is the past decelerating DM dominated solution and point $E_{0}$ where the universe ends in the accelerating completely DE dominated phase. Also note that, $E_{3},~{}E_{4}$ lie in the accelerating phase and on the other hand, at $E_{2}$, acceleration can not be determined. The graphics in the right plot of Fig. 10 properly represents the stability character of the critical points, leading to the results which are very same as the case $w<-1$ for the Models III.1 and III.4.
One-fluid system

### v2

@@ III.5 Model: $\Gamma=\Gamma_{0}H^{-2}$
In this class of matter creation models, a notable case is $\Gamma=\Gamma_{0}H^{-2}$, where $\Gamma_{0}$ is a constant with the same dimension as the cube of the Hubble parameter.
@@
where $\mu=\frac{\Gamma_{0}}{H_{0}^{3}}$ is a dimensionless parameter that adopts a positive value. To eliminate the singular behavior of the system (29) at $z=1$, we perform a time reparametrization via $dN=(1-z)^{3}d\tau$. This yields a reduced autonomous system, topologically equivalent to the original one, expressed as:
$\frac{dx}{d\tau}=$ $x(1-x)\left(3w(1-z)^{3}+\mu z^{3}\right),$ (30a)
$\frac{dz}{d\tau}=$ $\frac{3}{2}z(1-z)\left[\left(1+w(1-x)\right)(1-z)^{3}-\frac{\mu xz^{3}}{3}\right].$ (30b)
Here, one can easily conclude that $\bf R$} is a positively invariant set in the dynamical system (30) because the boundaries of $\bf R$} are invariant manifolds of (30). Here, $q=\frac{1}{2}\left[1+3w(1-x)-\frac{\mu xz^{3}}{(1-z)^{3}}\right]$. Therefore, at present time (i.e. $z=1/2$) for the accelerating expansion of the universe, the physical trajectories must follow the condition $(1-\mu)x+(1+3w)(1-x)<0$. The cosmological features that are obtained from the above system are briefly given in the Table 5. Now, we split the equation of the state parameter into various parts and analyze the system (30) in each part of $w$, which have been described below:
Critical point Existence Eigenvalue Stability $\mathbf{\Omega_{f}}$ $\mathbf{\Omega_{\rm dm}^{\rm eff}}$ $\mathbf{q}$ Acceleration
$E_{0}(0,0)$ Always $\left(3w,\frac{3}{2}(1+w)\right)$ Stable if w<1w<-1; 1 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Saddle if 1<w<0-1<w<0;
@@
$E_{2}(0,1)$ Always $(\mu,0)$ Non-hyperbolic Saddle if w>1w>-1; 1 0 Undefined Undetermined
Unstable if w<1w<-1
$E_{3}(1,1)$ Always $\left(-\mu,\frac{\mu}{2}\right)$ Always Saddle 0 1 -$\infty$ Yes
$E_{4}\left(1,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$ $\mu>0$ $\left(-\frac{\mu(1+w)}{\left(1+\sqrt[3]{\frac{\mu}{3}}\right)^{3}},-\frac{3\mu}{2\left(1+\sqrt[3]{\frac{\mu}{3}}\right)^{3}}\right)$ Stable if w>1w>-1 0 1 1-1 Yes
Saddle if w<1w<-1
$E_{5}\left(0,z_{c}\right)$ w=1w=-1 $\left(\mu z_{c}^{3}-3\left(1-z_{c}\right)^{3},0\right)$ Stable if $z_{c}<\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$ and 11 0 1-1 Yes
@@
$E_{6}\left(x_{c},\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$ w=1w=-1 $\left(0,-\frac{3\mu x_{c}}{2\left(1+\sqrt[3]{\frac{\mu}{3}}\right)^{3}}\right)$ Stable $1-x_{c}$ $x_{c}$ 1-1 Yes
@@
Table 5: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (30) with the matter creation rate $\Gamma=\Gamma_{0}H^{-2}$.
@@
Figure 13: Description of the phase space structure for the matter creation rate $\Gamma=\Gamma_{0}H^{-2}$. Upper Left Plot: The phase plot of the system (30) when we have assumed $w=0.1$ and $\mu=2$. For other values of $w>0$ and $\mu>0$, we can also obtain similar phase space structure. Upper Middle Plot: The phase space of the system (30) when the EoS $w$ takes the value 0. Here we use $\mu=2$ but any positive value of $\mu$ gives same type of phase portrait. Upper Right Plot: The phase plot of the system (30) considering w=0.3w=-0.3 and $\mu=3$. Also, we can get similar type of graphics for any positive value of $\mu$ and negative value of $w$ in the interval (1,0)(-1,0). Lower Left Plot: The phase space of the system (30) when the EoS $w$ takes the value 1-1. Here we use $\mu=2$ but any positive value of $\mu$ gives similar type of phase portrait. Lower Right Plot: The phase plot of the system (30) when we assume w=2w=-2 and $\mu=2$. For other values of w<1w<-1 and $\mu>0$, we can also obtain similar graphics. Here the green region corresponds to the decelerating phase ($q>0$), the cyan region represents the accelerating phase with 1<q<0-1<q<0 and the yellow region corresponds to the super accelerating phase (i.e. q<1q<-1). The black dotted curve separating the cyan and yellow regions (in the lower left graph this curve is not clear because of the critical line) corresponds to q=1q=-1.
Figure 14: We have presented the evolution of key cosmological parameters namely, the DM density parameter $\Omega_{\rm dm}^{\rm eff}$, the second fluid density parameter $\Omega_{f}$, the decelerating parameter $q$, and the effective equation of state of DM $w_{\rm dm}^{\rm eff}$ for the dynamical system (30) with the matter creation rate $\Gamma=\Gamma_{0}H^{-2}$. Upper Left Plot: This figure is drawn with the model parameters $w=0.1$, $\mu=1$, and it shows that the final fate of the universe is late time accelerated evolution in cosmological constant era dominated by DM. Upper Right Plot: This figure is produced for the parameters values $w=0$ and $\mu=1$. Here also, the final fate of the universe is late time accelerated evolution in cosmological constant era dominated by DM. Lower Left Plot: We have used w=1w=-1 and $\mu=1$. This plot depicts that the ultimate fate of the universe is found to be late time accelerated evolution in cosmological constant epoch which is achieved after the occurrence of a decelerated DM dominated phase. Lower Right Plot: We have taken w=1.2w=-1.2 and $\mu=1$. This plot highlights that the ultimate fate of the universe is found to be late time DE dominated accelerated evolution in phantom epoch which is achieved after the occurrence of a decelerated DM dominated phase.
[…]
@@
When $w>0$, the dynamical system (30) allows five critical points $E_{0},~E_{1},~E_{2},~E_{3}$ and $E_{4}$. Here, the accelerating phase includes the points $E_{3},~E_{4}$ and the decelerating phase contains the critical points $E_{0},~E_{1}$. Again, $E_{1},~E_{3},~E_{4}$ correspond to matter dominated solutions, while $E_{0},~E_{2}$ represent completely second fluid dominated points. One can easily conclude about stability from the direction of the vector field on the boundary of the phase space and signs of the eigenvalues. Therefore, $E_{0}$ is unstable, $E_{4}$ is the only globally stable point and $E_{1},~E_{2},~E_{3}$ are saddle by nature. Note that, $E_{4}$ is the DM dominated late time accelerating globally stable point. The upper left plot of Fig. 13 gives clear picture of the qualitative behavior of critical points. The inherent features of this plot are similar to the case $w>0$ of the Models III.1 and III.4. As further evidence, the upper left plot of Fig. 14 displays the evolution of the DM and second fluid density parameters $\left(\Omega_{\rm dm}^{\rm eff},\Omega_{f}\right)$ along with the deceleration parameter $q$ and the effective EoS of DM, $w_{\rm dm}^{\rm eff}$. The evolution of these parameters is almost similar to what we have observed earlier in the models with $\Gamma=\Gamma_{0}$, $\Gamma=\Gamma_{0}H^{-1}$.
@@
When $w=0$, we obtain three isolated critical points $E_{2}(0,1),~E_{3}(1,1)$ and $E_{4}\left(1,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$ along with one critical lines, namely, $E_{7}(x_{c},0)$. Now $x$ is always increasing and, $z$ is increasing and decreasing depending on $z$ variable chosen from $\left(0,\frac{1}{1+\sqrt[3]{\frac{\mu x}{3}}}\right)$ and $\left(\frac{1}{1+\sqrt[3]{\frac{\mu x}{3}}},1\right)$ respectively. Therefore $E_{2},~E_{3}$ are saddle, $E_{7}$ is unstable and $E_{4}$ is a globally stable critical point, which are obtained in the upper middle plot of Fig. 13. Once again $E_{4}$ corresponds to DM dominated late time accelerated globally stable point. In this case, the cosmological characteristics are completely identical with the case $w=0$ of the Models III.1 and III.4. Supporting this scenario, Fig. 14(upper right plot) illustrates the evolution of $\Omega_{\rm dm}^{\rm eff}$, $\Omega_{f}$, decelerating parameter $q$ and $w_{\rm dm}^{\rm eff}$. We also noticed that the evolution of these parameters is almost similar to what we have observed earlier in the models with $\Gamma=\Gamma_{0}$, $\Gamma=\Gamma_{0}H^{-1}$.
@@
When $\frac{1}{3}\leq w<0$, similar to the case where $w>0$, in this part of $w$, we acquire five isolated critical points which are depicted in the Table 5. From this table and the flow on the unit square, we can say that $E_{1}$ is unstable i.e. the DM dominated point $E_{1}$ leaves its decelerating phase, $E_{0},~E_{2},~E_{3}$ are saddle by behavior and once again, $E_{4}$ is the DM dominated accelerating globally stable point. Therefore, the universe ends with DM dominated accelerating phase. The orbits in the phase space are highlighted in the upper right plot of Fig. 13, providing the results which replicate the case 1/3<w<0-1/3<w<0 for the Models III.1 and III.4.
@@
When $1<w<-\frac{1}{3}$, this part of $w$ is analogous to the above case where $w$ lies in the interval $\left[-\frac{1}{3},0\right)$, except decelerating parameter at the critical point $E_{0}$ which belongs to the accelerating region in the phase space. Therefore, the phase plot is similar to the upper right plot of Fig. 13.
@@
When w=1w=-1, we obtain two isolated critical points $E_{1},~E_{3}$ which are mentioned in Table 5 and two critical lines, namely, $E_{5}(0,z_{c}),~E_{6}\left(x_{c},\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$. It is clear that $x^{\prime}$ is positive and negative respectively if we choose $z$ from the interval $\left(\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}},1\right]$ and the interval $\left[0,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$. Again $z$ is increasing and decreasing below and above the line $z=\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$ respectively. In the phase space below the line $z=\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$, the equation of the separatrix is $z=\sqrt[3]{\frac{3}{\mu}}(1-z)\sqrt{1-x}$ which connects the critical points $E_{1}$ and $\left(0,\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}\right)$. As a result, $E_{1}$ is unstable, $E_{3}$ is saddle, $E_{6}$ is stable and, $E_{5}$ is stable when $z_{c}<\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$ and unstable when $z_{c}>\frac{1}{1+\sqrt[3]{\frac{\mu}{3}}}$. The lower left plot of Fig. 13 depicts qualitative features of the critical points. Here, $E_{5}$, lying in the accelerating phase, corresponds to completely DE domination and $E_{6}$, belonging also to the accelerating phase, represents the points where DM, DE coexist except the end points. Clearly, $E_{6}$ can solve cosmic coincidence problem. The key findings of this case are equivalent to the Models III.1 and III.4 with w=1w=-1. Fig. 14 (lower left plot) offers further support, capturing the evolving dynamics of the DM and DE density parameters, along with the deceleration parameter $q$ and $w_{\rm dm}^{\rm eff}$.
@@
When w<1w<-1, one can get two unstable critical points $\left(E_{1},E_{2}\right)$, two saddle type critical points $\left(E_{3},E_{4}\right)$ and $E_{0}$ which is the only globally stable critical point. From Table 5, it is clear that, $E_{1}$ is the past decelerating DM dominated solution and point $E_{0}$ where the universe ends in the accelerating completely DE dominated phase. Also note that, $E_{3},~E_{4}$ lie in the accelerating phase and on the other hand, at $E_{2}$, acceleration can not be determined. The graphics in the lower right plot of Fig. 13 properly represents the stability character of the critical points, leading to the results which are very same as the case w<1w<-1 for the Models III.1 and III.4. Additional insight is provided by the lower left plot of Fig. 14, where the evolution of $\Omega_{\rm dm}^{\rm eff}$, $\Omega_{f}$, $q$ and $w_{\rm dm}^{\rm eff}$ is clearly exhibited, further substantiating the model’s predictions.
Finally, in Fig. 15 we compare the normalized Hubble parameter $H/H_{0}$ for the matter creation rate ($\Gamma=\Gamma H^{-2}$) with the standard $\Lambda{\rm CDM}$ model (upper panel). The lower panel shows $\Delta H/H$. The evolution of the matter creation model follows $\Lambda{\rm CDM}$ at low redshift but differs at higher redshifts due to particle creation effects.
III.6 Model: $\Gamma=2\left(1-\frac{1}{\xi}\right)\frac{\dot{H}}{H}$
In this section, we consider an inhomogeneous matter creation rate given by $\Gamma=2\left(1-\frac{1}{\xi}\right)\frac{\dot{H}}{H}$, where $\xi$ is a non-zero constant. We then analyze how this specific form of $\Gamma$ influences the evolution of the Universe.
The autonomous dynamical system, formulated using the dimensionless variables $x$ and $z$ introduced in eqn. (20), can be read as
$x^{\prime}=$ $\frac{3x(1-x)(1+w-\xi)}{\xi(1-x)+x},$ (31a)
$z^{\prime}=$ $\frac{3\xi z(1-z)(1+w(1-x))}{2(\xi(1-x)+x)}.$ (31b)
The autonomous system described above may exhibit singular behavior depending on the value of the parameter $\xi$. For positive values of $\xi$, the system remains free of singularities. However, when $\xi$ is negative, singularities arise along the line $x=\frac{\xi}{\xi-1}$. For this matter creation rate, the decelerating parameter is given by $q=\frac{\xi+2(\xi-1)x+3w(1-x)}{2(\xi-(\xi-1)x)}$. We now execute a phase space stability analysis by considering two distinct regimes of the model parameter $\xi$: the positive $\xi$ region and the negative $\xi$ region. The critical points or lines, along with the corresponding values of key cosmological parameters, are summarized in Table. 6.